OpenStudy (anonymous):

How much work is required to lift an object with a mas of 5.0 kilograms to a height of 3.5 meters?

OpenStudy (ghazi):

work done = potential energy stored = mgh = 5*10*3.6= 180N

OpenStudy (anonymous):

But the answer chooses are A.17 joules B.1.7 × 10^2 joules C.1.9 × 10^2 joules D.2.8 × 10^2 joules

OpenStudy (ghazi):

well i guess options are incorrect....or else you can convert 180 N into joules 1joule= 1 n-m

OpenStudy (anonymous):

Wow ok thank you I will go tell her this

OpenStudy (ghazi):

anytime

OpenStudy (anonymous):

1.76*10^2 joules

OpenStudy (anonymous):

work done = potential energy stored = mgh = 5*9.8*3.6= 11760N-m or 1.76*10^2joules @jrkiller

OpenStudy (anonymous):

@ghazi the unit for work is joules, not newtons. so ur ans will anyway be in J. @theyatin ur ans should be 176.4. J....not 11760 N-m. @jrkiller as theyatin said, the ans is opt B. there is a difference between theyatin and ghazi's answer, because both took different values of g. ghazi took g=10 and theyatin took g=9.8.

OpenStudy (ghazi):

yea i forgot to write N-m

OpenStudy (anonymous):

@Vaidehi09 its a type ddear i wanted to type 1760 just to round off dont be so rude to kids ;'(

OpenStudy (anonymous):

@theyatin did i sound rude?! sorry...no intention to!!

OpenStudy (anonymous):

lol bacha i was kidding :) just don't give so much stress on correcting others . . .

OpenStudy (anonymous):

@ghazi no offense to u too!! @theyatin right. i'll be careful from now on. thanks! :)

OpenStudy (ghazi):

@Vaidehi09 it's cool

OpenStudy (anonymous):

@ghazi :) thanks.